### v1

# Extracting Information in a Low-resource Setting: Case Study on Bioinformatics Workflows

Abstract
Bioinformatics workflows are essential for complex biological data analyses and are often described in scientific articles with source code in public repositories. Extracting detailed workflow information from articles can improve accessibility and reusability but is hindered by limited annotated corpora. To address this, we framed the problem as a low-resource extraction task and tested four strategies: 1) creating a tailored annotated corpus, 2) few-shot named-entity recognition (NER) with an autoregressive language model, 3) NER using masked language models with existing and new corpora, and 4) integrating workflow knowledge into NER models. Using BioToFlow, a new corpus of 52 articles annotated with 16 entities, a SciBERT-based NER model achieved a 70.4 F-measure, comparable to inter-annotator agreement. While knowledge integration improved performance for specific entities, it was less effective across the entire information schema. Our results demonstrate that high-performance information extraction for bioinformatics workflows is achievable.
Keywords:
Natural Language Processing Information Extraction Bioinformatics Workflows
1 Introduction
In data-intensive sciences, such as bioinformatics, scientific results are produced using complex data analysis pipelines taking in vast amounts of experimental data. These pipelines consist of numerous steps making calls to bioinformatics tools and require considerable computational time. They can be implemented using scripts (Bash, Python, etc.) that orchestrate tool execution and serve as the glue between data, processes, tools, and the execution environment. However, the development and use of such scripts lead to challenges in the design and execution of the processing chain, as well as maintenance and reproducibility issues. In response to these challenges, efforts have been made over the past two decades to provide bioinformaticians with approaches that guide them toward better automation of data analysis pipelines: the scientific workflow systems [4]. Two systems stand out today in bioinformatics: Nextflow [5] and Snakemake [15].
Currently, we observe: (1) an increase in scientific articles describing bioinformatics workflows in a descriptive form (steps described without being executable) and (2) a rise in bioinformatics workflows available on repositories like GitHub that are in a programmatic form (implementation available but the lack of documentation makes them difficult to reuse).
In natural language processing (NLP), bioinformatics workflows remain underrepresented. To our knowledge, there is no comprehensive corpus encompassing a rich and diverse set of entities describing workflows (bioinformatics tools, data, management systems, etc.). However, some corpora focus on specific types of entities, such as those centered on bioinformatics tools [6, 13]. This lack of resources makes bioinformatics a low-resource context, making it difficult to structure information or analyze texts with the limited annotated data available.
To overcome the low-resource challenges in the field, several strategies can be employed. First, decoder-based approaches, such as those using generative models, offer a potential solution, although they may not always outperform other methods in terms of extraction accuracy [3, 17]. Next, encoder-based approaches can be explored in two ways: using larger, related corpora to enrich model training, or by developing a small, domain-specific training corpus to focus on the extraction of entities relevant to the field. Finally, another promising approach is the injection of domain-specific knowledge directly into language models.
The challenge is thus to develop a simple and automated approach capable of extracting information about workflows described in the literature, aiming to not only systematically document workflows presented in the literature but also facilitate their registration in workflow repositories. This publication introduces a methodology for extracting information about workflows from the literature.
The contributions of this work are as follows: (a) a representation of workflow components using a schema comprising 16 entities and a corpus manually annotated named BioToFlow; (b) experiments with few shot Named-Entity Recognition (NER) using an autoregressive language model; (c) NER experiments using a pre-existing corpus and BioToFlow and (d) integration of workflow knowledge into NER models.
2 Workflow Information Representation
2.1 Workflow Representation Schema
To systematically describe the bioinformatics workflows and their components, we relied on discussions with workflow experts (notably from the ELIXIR and BOSC International Bioinformatics communities) and the review of several articles describing workflows (e.g., [14, 25, 4, 5, 15]). Generally, bioinformatics workflows consist of data analysis steps, each containing a script that may or may not invoke a bioinformatics tool. For a workflow to be executed, it is crucial to maintain a record of its execution environment. Figure 1 presents the proposed schema, which distinguishes between three main categories of entities. The first category is core entities, representing major components of a workflow, such as bioinformatics tools and data. Tools in this category can also be further classified based on their context: whether they are explicitly identified as bioinformatics tools (BioInfo), laboratory-related tools (Lab), tools mentioned but not used in the final workflow (Context), or tools referenced in a general sense without specific details (General). The second category includes environment entities, which capture the resources and configurations required for workflow execution, such as the execution environment or programming languages used in scripts. Finally, the third category addresses more specific details, including versioning information, bibliographic references describing tools or libraries, as well as descriptions and parameters associated with certain elements.
Figure 1: Bioinformatics workflow representation schema.
2.2 Annotation of Workflow Information in Articles: BioToFlow
2.2.1 Article Selection Criteria
Articles describing bioinformatics workflows with a link to workflow code in Github have been targeted using a query on PubMed and PubMedCentral [19] [^1] PubMed query: (nextflow[Abstract] OR snakemake[Abstract] OR nextflow[Title] OR snakemake[Title]) AND github[All Fields]. As of 11/18/2024, it returns 124 and 123 articles describing workflows from Nextflow and Snakemake systems respectively.
2.2.2 Creation of an Annotated Corpus
To validate this information representation schema and create a resource for evaluating the extraction of such information, we developed an annotated corpus using a collaborative method [7] with the help of an annotation guide that we designed. Annotations were performed using BRAT (Brat Rapid Annotation Tool [21], v1.3 p1), recognized for its qualities including the ability to process nested entities, an existing annotation schema, and pre-annotations [18].
Seven annotators participated in the annotation campaigns, which were conducted in multiple stages: the annotators first worked on common texts, then on different texts. Inter-annotator agreement (IAA) was calculated using the F-measure [^2] For a detailed review of inter-annotator agreement measures and their usage context, we refer the reader to [1]. In the case of named entity annotation, [9] demonstrates that the F-measure and $\kappa$ can be considered equivalent. with BRAT-Eval [23](v0.0.2), a tool that compares two sets of annotations and analyzes divergences. The evaluation of agreement can be strict (agreement between two annotations only if identical text portions are annotated with the same label) or relaxed (agreement if two identical or overlapping text portions are annotated with the same label). The score ranges from 0 (no agreement) to 1 (perfect agreement). IAA obtained were all higher than 70% in relaxed mode. When we look in more detail at the extraction scores for each entity, we found a heterogeneity in the scores showing that some are easier to annotate than others. There are several causes for these differences, including ambiguity in the tagset, criteria for discrimination [8], and, finally, the expertise domain of the annotators. Figure 2 shows entity annotations using our schema on an excerpt from an article describing a Nextflow workflow (PMID 35171290).
Figure 2: Excerpt from the annotated corpus using the BRAT software.
Overall, our corpus, named BioToFlow [^3] https://doi.org/10.5281/zenodo.11204427, is a small corpus composed of 52 articles (26 articles related to Nextflow workflows and 26 on Snakemake workflows) with a total of 78,419 tokens (27,786 annotated tokens). However, the entities are diversified and distributed according to Table 1. In our corpus, we allow the presence of nested entities ($\sim$8of entities in BioToFlow; an example can be seen in Figure 2 where a WorkflowName entity is nested within a Biblio entity).
Table 1: BioToFlow: number of occurrences per entity.
Entities Occurrences Entities Occurrences
Data 2,434 Version 454
Tool 1,482 Hardware 429
Description 1,300 Database 288
Biblio 1,251 ManagementSystem 243
Method 936 Container 108
WorkflowName 851 ProgrammingLanguage 104
File 780 LibraryPackage 101
Parameter 464 Environment 83
Using this corpus and its annotations as a starting point, we explored different extraction approaches, starting with decoder-based approaches and then moving on to encoder-based approaches.
3 Decoder-based Approach for Named Entity Recognition
Because of the small size of the BioToFlow corpus, we sought to leverage an auto-regressive language model to perform few-shot named entity recognition. Specifically, we used Llama-3-8B-Instruct, which demonstrated the best performance among generative models in a comparative study of masked language models and generative models utilizing prompting techniques [17]. Following the same methodology as in this study, we conducted a series of few-shot experiments, varying the number of examples provided and testing different prompt formulations. We use random sampling to split the BioToFlow corpus: we use 75% of the corpus (39 articles) for training and the remaining 25% (13 articles) for testing. The results obtained on the BioToFlow test set using the Llama-3-8B-Instruct model with the best hyperparameters are presented in Table 2.
Table 2: Performance of Llama-3-8B-Instruct model with the best hyperparameters (relaxed mode).
Entities P R F1 Entities P R F1
Biblio 93.1 82.2 87.3 LibraryPackage 3.8 5.6 4.5
Container 29.2 50.0 36.8 ManagementSystem 41.1 38.3 39.7
Data 31.2 27.6 29.3 Method 15.4 12.1 13.6
Database 28.3 24.6 26.3 Parameter 11.4 8.8 9.9
Description 26.2 19.4 22.3 ProgrammingLanguage 23.5 14.3 17.8
Environment 60.0 13.0 21.4 Tool 54.9 59.1 56.9
File 30.7 34.8 32.6 Version 67.1 39.5 49.7
Hardware 20.5 26.3 23.1 WorkflowName 29.1 37.7 32.8
Overall 40.1 36.7 38.3
We obtained an overall performance below 40% in relaxed mode and 31.2% in strict mode (not shown). These results are consistent with findings on other biomedical corpora [17] and lead us to explore alternative methods leveraging encoder-based models.
4 Encoder-based Approach for Named Entity Recognition
Encoder-based approaches for NER typically require a substantial amount of annotated corpus. In our experiments, we leverage a large corpus with annotations covering a subset of workflow information and a small corpus with comprehensive coverage of workflow information.
4.1 Experiments on a Large Corpus: SoftCite
To our knowledge, no corpus fully covers workflows. However, we identified corpora with annotations partially aligned with our schema, such as those focusing on datasets [20] or bioinformatics tools. We focused on bioinformatics tools, as identifying them offers crucial insights into a workflow’s functionality. We can cite the CZ Software Mentions dataset [13], which includes software mentions extracted using a trained SciBERT model from diverse sources like PubMed Central and the Chan Zuckerberg Initiative. Another relevant corpus is SoftCite [6], a gold-standard dataset where all articles were manually annotated. SoftCite comprises 4,971 articles from the fields of biomedicine and economics. We chose SoftCite due to its manual annotation process and, to assess its relevance, we selected a subset of 1,159 articles specifically related to biology.
Since SoftCite does not contain the same entity types as BioToFlow, our first step was to identify possible overlaps between the two entity schemas and perform a mapping to convert SoftCite entity names to those used in BioToFlow. Table 3 represents the correspondence between entities in SoftCite (left column) and those in BioToFlow (right column). Some entities in the SoftCite corpus, publisher_person, figure, table, and formula, have no equivalent in BioToFlow. Similarly, certain entities in BioToFlow, including data, container, hardware, and description, are not present in SoftCite. After the conversion, the number of entities and occurrences are presented in Table 4.
Table 3: Entity correspondence between SoftCite and BioToFlow.
SoftCite BioToFlow
software Tool
environment Tool
url Biblio
component LibraryPackage
implicit Tool_general
publisher Biblio
environment Environment
bibr Biblio
version Version
url Biblio
language ProgrammingLanguage
Table 4: SoftCite: Number of occurrences per entity.
Entities Occurrences Entities Occurrences
Tool 4,142 Biblio 3,946
Version 1,349 LibraryPackage 65
ProgrammingLanguage 61
For the extraction of our target entities, we chose a biLSTM-CRF neural model [24], implemented by the NLStruct Python library package [^4] https://github.com/percevalw/NLStruct (v0.2.0). This library offers the dual advantage of supporting the detection of nested entities, which is essential for certain entity types in our dataset, and accepting input files in the BRAT format. Additionally, it has demonstrated strong performance in the biomedical domain, which closely aligns with our field. More specifically, we used this model with the SciBERT language model [2], more specifically trained on scientific articles. We conventionally chose to randomly divide the articles into two main sets: the first set, 75% of the corpus (927 articles), was used to create the training set, while the remaining 25% (232 articles) constituted the test set. Within the training set, two-thirds of the articles (649 articles) were used for the actual training of the models and one-third (278 articles) for their validation. A distinctive aspect of our approach is that we constructed five splits through random sampling to minimize dependency on a specific division. For each split, five versions of the entity recognition model were produced using different random seeds, ensuring more robust and reproducible results.
Table 5: Average performance metrics on five splits and five random seeds (with standard deviation) of percentage scores obtained using SciBERT and NLStruct in relaxed mode on SoftCite corpora.
Softcite
P R F1
Biblio 90.6$\pm$1.4 99.0$\pm$0.4 94.6$\pm$0.9
Environment 0.0$\pm$0.0 0.0$\pm$0.0 0.0$\pm$0.0
LibraryPackage 2.5$\pm$2.4 14.3$\pm$10.5 4.1$\pm$3.9
ProgLanguage 48.9$\pm$5.3 88.8$\pm$4.8 62.8$\pm$4.8
Tool 73.4$\pm$1.9 51.0$\pm$0.6 60.2$\pm$1.0
Version 49.8$\pm$2.2 93.5$\pm$0.6 65.0$\pm$1.9
Overall-focused 73.1$\pm$1.1 70.7$\pm$0.6 71.9$\pm$0.8
We test the results on 13 test articles from BioToFlow to see if the entities of the SoftCite corpus could be used to train and extract a part of the BioToFlow entities. The performance on a subset of entities (Table 5) using this corpus is significantly better than those achieved with a decoder-based approach.
4.2 Experiments on a Small Corpus: BioToFlow
Similar to the SoftCite corpus experiments, we conducted experiments on BioToFlow using NLStruct and divided the data as follows: 75% of the corpus (39 articles) in the training set, with two-thirds of these (26 articles) used for model training and one-third (13 articles) for validation and 25% (13 articles) formed the test set. The results obtained are presented in Table 6.
Table 6: Average performance on five splits and five random seeds (with standard deviation) of percentage scores obtained with SciBERT and NLStruct in relaxed mode on the BioToFlow corpus. Overall-focused results are directly comparable to Table 5.
Entities P R F1 Entities P R F1
Biblio 97.9$\pm$0.4 97.4$\pm$1.0 97.7$\pm$0.4 LibraryPackage 39.8$\pm$12.3 30.9$\pm$6.7 33.9$\pm$7.4
Container 78.3$\pm$11.2 80.2$\pm$12.5 78.9$\pm$11.8 MgmtSystem 72.9$\pm$5.6 90.9$\pm$4.0 80.6$\pm$3.4
Data 70.9$\pm$2.1 57.1$\pm$2.3 63.2$\pm$1.9 Method 38.6$\pm$2.6 68.3$\pm$3.4 49.2$\pm$2.9
Database 63.8$\pm$3.6 54.8$\pm$4.3 58.7$\pm$1.8 Parameter 53.2$\pm$3.3 53.1$\pm$4.8 52.9$\pm$2.9
Description 65.4$\pm$3.9 59.3$\pm$1.8 62.1$\pm$1.5 ProgLanguage 86.9$\pm$2.1 87.2$\pm$2.9 86.9$\pm$2.4
Environment 50.3$\pm$17.9 79.4$\pm$10.1 60.3$\pm$17.3 Tool 77.7$\pm$3.2 72.2$\pm$2.8 74.8$\pm$1.0
File 84.3$\pm$0.5 82.5$\pm$1.4 83.4$\pm$0.6 Version 86.8$\pm$2.6 90.8$\pm$2.0 88.7$\pm$0.9
Hardware 65.8$\pm$1.2 79.2$\pm$4.4 71.7$\pm$1.7 WorkflowName 67.1$\pm$2.2 77.8$\pm$4.4 72.0$\pm$2.1
Overall 71.2$\pm$1.3 69.6$\pm$0.9 70.4$\pm$0.8 Overall-focused 85.5$\pm$1.2 83.4$\pm$1.2 84.4$\pm$0.5
The first results show that training a model on our little corpus obtains higher results for all the entities (see Table 2, 5, and 6).
4.3 Fusion of SoftCite and BioToFlow Corpora
The previous results show that SoftCite and BioToFlow, when used independently, achieve good scores (above 70%). We also merged the two corpora and trained models on the combined dataset to see the contribution of the SoftCite corpus. Two configurations of the SoftCite corpus were considered: (a) entities were only converted to match BioToFlow entity names and (b) each SoftCite article was annotated using predictions from a model trained on BioToFlow.
The scores for specific entities are provided in Table 7. The results suggest that the SoftCite corpus, in its silver form, achieves a slightly higher F1 score compared to a model trained only on BioToFlow (Table 6). These corpus fusions resulted in some entity scores increasing while others decreased.
Table 7: Average performance metrics on five splits and five random seeds (with standard deviation) of percentage scores obtained using SciBERT and NLStruct in relaxed mode on the fusion of BioToFlow and SoftCite corpora.
With SoftCite conversion With SoftCite silver
P R F1 P R F1
Data 56.9$\pm$2.3 64.1$\pm$1.5 60.1$\pm$0.7 75.8$\pm$2.6 56.8$\pm$2.1 64.9$\pm$0.8
Description 62.9$\pm$3.7 62.5$\pm$1.3 62.6$\pm$1.2 71.0$\pm$0.8 57.4$\pm$0.9 63.5$\pm$0.7
LibraryPackage 36.2$\pm$7.9 38.1$\pm$6.3 36.1$\pm$7.0 25.3$\pm$6.4 37.2$\pm$8.9 29.0$\pm$5.7
ProgLanguage 88.6$\pm$5.4 87.6$\pm$4.0 87.9$\pm$2.4 90.3$\pm$1.3 88.3$\pm$3.0 89.2$\pm$1.4
Tool 76.9$\pm$1.3 72.2$\pm$1.2 74.5$\pm$0.9 78.8$\pm$1.9 70.0$\pm$0.9 74.1$\pm$1.1
Overall 64.4$\pm$1.0 73.6$\pm$0.9 68.7$\pm$0.3 73.0$\pm$0.5 68.5$\pm$1.4 70.7$\pm$0.6
5 Integration of Knowledge into Language Models
Since the various tested approaches did not show any significant improvement in NER scores compared to a masked language model trained on BioToFlow, we explored one final approach: the integration of knowledge into language models.
5.0.1 Used of a Pre-initialized Encoder
In the spirit of work exploiting pretraining for NER [16], we trained models on BioToFlow following the same methodology as in Section 4.2 but with an encoder initialized with one of the models trained during the NER experiments on the SoftCite corpus (see Section 4.1). The scores of Table 8 for a subset of entities show that the results obtained are nearly identical to those achieved with a model whose encoder was not pre-initialized.
Table 8: Average performance metrics on five splits and five random seeds (with standard deviation) of percentage scores for a subset of entities obtained using SciBERT and NLStruct in relaxed mode on BioToFlow corpus with the encoder initialization.
P R F1 P R F1
Database 63.9$\pm$3.2 56.2$\pm$3.3 59.7$\pm$1.3 Method 38.2$\pm$1.2 68.1$\pm$3.3 48.9$\pm$1.6
Environment 47.7$\pm$18.9 80.0$\pm$14.3 58.5$\pm$20.1 ProgLanguage 87.1$\pm$1.0 85.9$\pm$5.0 86.2$\pm$2.5
LibraryPackage 40.2$\pm$11.5 32.4$\pm$5.6 35.1$\pm$5.9 Tool 77.3$\pm$3.1 72.1$\pm$2.2 74.5$\pm$0.6
MgmtSystem 72.7$\pm$5.9 94.4$\pm$5.7 82.0$\pm$4.2 Overall 71.3$\pm$0.8 69.6$\pm$1.3 70.4$\pm$0.8
5.0.2 Adding Vocabulary
To improve the extraction of bioinformatics tool names, another strategy involves incorporating external knowledge into the encoder. Since the knowledge to inject in our case is made of lists of names, we chose to perform this injection at the level of its vocabulary, similarly to work such as [11]. Various knowledge bases of bioinformatics tools exist, from which tool and binary names can be extracted. We utilized four knowledge bases: Biotools [12], Bioconda [10], Biocontainers [22] and Bioweb [^5] https://bioweb.pasteur.fr/welcome. From these knowledge bases, we aggregated the names of tools and binaries to create an enriched vocabulary. This new vocabulary was then injected into the SciBERT language model to improve its ability to recognize bioinformatics tool names. We evaluated this strategy by conducting two experiments on BioToFlow, testing two approaches: (a) fine-tuning the SciBERT model with the new vocabulary and (b) using the enriched vocabulary without additional fine-tuning. In both cases, the embeddings of the added tokens were initialized by averaging the embeddings of their subtokens.
Adding bioinformatics tool vocabulary appears to improve the extraction of tool names, particularly when fine-tuning the SciBERT model after incorporating the new vocabulary, with an F1 score increasing from 74.8$\pm$1.0to 77.0$\pm$2.8%. As observed in previous experiments, some entities benefit from the added knowledge, while others experience slight decreases in performance (see Table 9). Specifically, the Data entity shows improvements with vocabulary addition in both fine-tuning and non-fine-tuning scenarios. Entities like LibraryPackage and Tool also benefit from fine-tuning with the added vocabulary. Conversely, scores for ManagementSystem and ProgrammingLanguage entities show slight declines.
Table 9: Average performance metrics on five splits and five random seeds (with standard deviation) of percentage scores obtained using SciBERT and NLStruct in relaxed mode with the addition of bioinformatics tool vocabulary, with or without fine-tuning.
Add voc – finetune Add voc – no finetune
P R F1 P R F1
Data 70.0$\pm$7.2 63.4$\pm$10.9 66.4$\pm$8.9 71.7$\pm$2.4 56.9$\pm$2.8 63.4$\pm$2.3
LibraryPackage 52.5$\pm$7.1 57.4$\pm$15.6 52.3$\pm$10.0 37.1$\pm$8.0 30.4$\pm$7.2 32.7$\pm$6.0
ManagementSystem 75.6$\pm$7.9 76.8$\pm$6.7 75.5$\pm$7.0 69.7$\pm$3.9 93.4$\pm$3.3 79.7$\pm$1.7
ProgrammingLanguage 81.9$\pm$12.6 82.1$\pm$11.5 81.9$\pm$12.0 86.0$\pm$2.1 85.8$\pm$2.3 85.8$\pm$2.1
Tool 77.2$\pm$3.7 76.9$\pm$2.7 77.0$\pm$2.8 74.5$\pm$2.5 73.9$\pm$2.6 74.1$\pm$0.3
Overall 70.4$\pm$1.7 70.2$\pm$1.5 70.3$\pm$0.4 70.6$\pm$1.2 69.5$\pm$1.2 70.1$\pm$0.9
6 Conclusion
Based on a workflow representation schema, we developed a small entity-rich corpus, BioToFlow. Using this corpus, we experimented with various entity extraction techniques, including decoder-based models, which achieved scores below 50%, and encoder-based methods. These included leveraging a large corpus with annotations covering a subset of the schema and our BioToFlow corpus. A final approach sought to incorporate additional knowledge into the models. The most effective strategy involved training a model on our small corpus while integrating a vocabulary list specific to bioinformatics tools, which improved the extraction of this entity type. To extend this latest approach, we will investigate various ways to enhance the embeddings of the added vocabulary, starting with their adaptation during the training of the NER models, similarly to [11]. More globally, the entity linking between the components extracted from workflows described in scientific literature and those found in code from public repositories is an essential task that we will consider to merge information from both sources.
{credits}
6.0.1 Acknowledgements
We sincerely thank N. Bossut, A. Gaignard, G. Marchment, M. Schmit, H. Menager, F. Lemoine for their help with the annotation process. This work has received support from the French government (Agence Nationale pour la Recherche) under the France 2030 program grant agreement ANR-22-PESN-0007 (ShareFAIR).
[…]
The authors have no competing interests.